OpenStudy (anonymous):

Correct the following samples of gases to standard temperature. (Assume pressure is constant) 1. 1622 cm3 at 15.0 oC

OpenStudy (stmoney):

Standard temp is 0 oC, so you will need to change temp to 0 oC. This means you will need to use charles's law. Temp must also be in Kelvin. V1/T1 = V2/T2. 1622 cm3/(15+273)=x/273. The volume is 1537.52cm3